what is 4 3/8 X 6 2/5 ( I forgot the rules )
what is ( adding ) 17 5/6 and 45/4 ( forgot the rules )
Thanks in advance for you assistance

Answer by 2139(3)

4 3/8 x 6 2/5
4 times by 8 (the denominator) and add by 3
6 times by 5 (the denominator) and add by 2
which is:
4x8= 24 + 3= 27/8
6x5= 30 + 2= 32/5
then find the LCM of 8 and 5= 40 (the denominator)
hence; 27(5)/8(5)= 135/ 40 = 3.375
32(8)/5(8)= 256/ 40 = 6.4
therefore; 3.375 x 6.4 = 21.6